# Intention Analysis Makes LLMs A Good Jailbreak Defender

### v2

@@ 4.2 Main Results
Our method significantly boosts safety against varying jailbreak attacks.
The performance of our method on SAP200 and DAN datasets is presented in Table 1&2, respectively. We can observe that our $\mathbb{IA}$ consistently achieves marked safety improvements against the most advanced and widespread jailbreak prompt attacks, across model sizes and alignment levels. Particularly for SAP200, the substantial reductions of ASR for Vicuna-7B (73.09%) and Vicuna-13B (70.38%) demonstrate the effectiveness. As for different alignment levels, not only does our method show satisfactory performance on SFT models, i.e. Vicuna-7B and 13B, but such enhanced safety also extends to well-aligned models like GPT-3.5, which registers nearly 0% ASR under our $\mathbb{IA}$.
Figure 4: The confusion matrix illustrating the relationship between the success of intention analysis and the harmlessness of the response. “IR Succ.” and “IR Fail.” are success or failure of intention analysis, respectively.
@@
Our method can consistently enhance model safety in the context of multilingual jailbreak.
Recent studies Deng et al. (2023b); Yong et al. (2023) reveal that the multilingual jailbreak poses a new defence challenge for LLMs. To verify the effectiveness of our method in multilingual jailbreak scenarios, we follow Deng et al. (2023b) to reproduce their settings on nine different languages with two jailbreak scenarios, including Chinese (zh), Italic (it), Vietnamese (vi), Arabic (ar), Korean (ko), Thai (th), Bengali (bn), Swahili (sw), and Javanese (jv). The results of vanilla ChatGPT and our $\mathbb{IA}$ are presented in Figure 3. We show that 1) our $\mathbb{IA}$ can consistently maintain performance, even in low-resource languages, e.g., th, bn, sw, and jv, 2) even in scenarios where a malicious jailbreak prompt [^6] We adopt the same jailbreak prompt in Deng et al. (2023b), namely AIM. is attached to the multilingual attacks, our $\mathbb{IA}$ still manages to reduce the ASR from around 80% to almost 0%. This demonstrates the effectiveness of our method under multilingual jailbreak attacks.
Our method maintains the model’s helpfulness on general tasks.
@@
To explore the impact of our method on the general performance of LLMs, we conduct experiments on several acknowledged helpfulness datasets and report the results in Table 4. As observed, our $\mathbb{IA}$ does not compromise the general performance, indicated by a comparable win rate on AlpacaEval and a similar accuracy on MMLU. Furthermore, $\mathbb{IA}$ can improve the truthfulness over different models in terms of better MC1 and MC2 scores. These results indicate that $\mathbb{IA}$ can be deployed in real applications to enhance LLM safety while preserving general helpfulness.
Our method consistently outperforms other defense baselines.
Table 3 lists the comparison results between our $\mathbb{IA}$ and the baselines. [^7] Due to computational resource and financial limitations, we randomly select 320 samples each from DAN and SAP200 datasets for comparative experiments. As observed, our method consistently shows superiority over other baselines on different datasets and model scales. Specifically, our $\mathbb{IA}$ outperforms the second-best method by 30.32% and 23.77% averagely on SAP200 and DAN, respectively. In addition, although ICD and System-Mode Self-Reminder achieve considerable reduction in ASR on AdvBench, their performance severely degrades when dealing with complex and stealthy jailbreak prompts. On the contrary, our method consistently outperforms other baselines across both prompt-level and token-level jailbreak datasets. We conjecture such satisfactory performance to be attributed to our design in intention analysis mechanism, which leverages LLM’s awareness of harmful content to enhance safety against jailbreak attacks.
@@ 5 Discussion
5.1 Analysis of Major Components in $\mathbb{IA}$
@@
LLMs can indeed recognize the intentions behind jailbreak queries.
To verify this, we count the samples where intentions are successfully recognized and further examine the correlation between recognizing intentions and producing harmless responses, as shown in Figure 4. We show that LLMs are highly effective in analyzing intentions of complex jailbreak queries, achieving a 100% success rate in most models, such as Vicuna-7B, Vicuna-13B, and MPT-30B-Chat. Moreover, the notable volume of harmless responses following successful intention recognition demonstrates the critical role of intention analysis in enhancing LLM safety.
Our method achieves comparable safety of well safety-trained LLMs while surpassing them in helpfulness.
Our method aims to enhance LLM safety in the inference stage. A natural question arises how does its performance compare to well safety-trained LLMs? To answer this, we compare our method with a representative well safety-trained LLM, i.e., LLaMA2-7B-Chat. The related results are listed in Table 7. We can observe that our method achieves comparable performance to LLaMA2-7B-Chat on safety datasets while significantly outperforming LLaMA2-7B-Chat on the helpfulness dataset by almost 5%. This demonstrates the advantage of our $\mathbb{IA}$ to achieve both safety and helpfulness goals.
Our method still works well if combining two stages into one.
As aforementioned, to maximize the performance, our $\mathbb{IA}$ follows a two-stage process. A natural question arises of whether our mechanism can be merged into one step, to save the decoding overhead. To verify this, we design a cheap one-pass $\mathbb{IA}$, without letting the LLMs separately explain the intention. As shown in Table 5, we see that while two-stage $\mathbb{IA}$ can achieve a relatively lower ASR, the one-pass $\mathbb{IA}$ also performs well, regarding the significantly ASR reduction of 61.7% and 64.27% for Vicuna-7B and Vicuna-13B, respectively.
Vicuna-7B Vicuna-13B ChatGLM-6B
Vanilla 67.2 65.4 45.8
+ One-Pass $\mathbb{IA}$ 5.50 1.13 39.0
+ Two-Stage $\mathbb{IA}$ 0.31 1.12 6.12
Table 5: Comparison of our $\mathbb{IA}$ with different implementations (one-pass and two stage) on SAP200 in terms of ASR. The best results are highlighted in bold.
Vicuna-7B Vicuna-13B ChatGLM-6B
Vanilla 67.2 65.4 45.8
+ Prompt A 2.94 0.88 5.81
+ Prompt B 5.13 2.06 5.44
+ Ours 0.31 1.12 6.12
Table 6: Ablation of different intention analysis prompts on SAP200 in ASR (%). The best and second best results are highlighted in bold and underline.
5.2 Robustness of Our $\mathbb{IA}$ Mechanism
Our method is robust to different prompts.
To assess the resilience of our method against specific expressions, we construct other two sets of alternative instructions for intention analysis. Comparison results are listed in Table 6. We show that these three prompt sets all result in significant and comparable decreases in ASR on the SAP200 across different models. This uniformity highlights a key conclusion: it is the intention analysis mechanism itself, but not the specific expression of the instructions, that makes our method effective.
Our method is robust against varying intention analysis quality.
A crucial question that emerges regarding our method is whether its performance would be significantly influenced by the quality of intention analysis. To address this concern, we simulate varying correct intention ratios (representing intention analysis quality) in the first stage by substituting a proportion of the intention with “[secret]”, and the corresponding performance of our method is presented in Figure 5. Key observations include 1) Under different correct intention ratios, the ASR values achieved by our method display a relatively stable trend (with variations within 10%), particularly for larger models (e.g., Vicuna 13B), indicating the robustness of our method to varying intention analysis quality, 2) the performance of our method gradually improves as the correct intention ratios increase, indicating that enhancing the ability to analyze intentions can further boost the effectiveness of our $\mathbb{IA}$.
@@ 5.2 Robustness of Our $\mathbb{IA}$ Mechanism
Figure 5: Performance of our method on DAN with varying correct intention ratio, which demonstrates the robustness to intention analysis ratio of our $\mathbb{IA}$.
@@
Table 7: Comparison between our method and well safety-trained LLM in safety and helpfulness. The best&second best are highlighted in bold and underline.
5.3 Human Evaluation
To validate GPT-3.5 as the harmfulness annotation function, we conduct a human validation process. First, we randomly select 200 responses from both the vanilla model and our method on the DAN dataset. Then we engage two expert annotators proficient in alignment studies of LLMs and fluent in English. After reassessing their evaluations for cases with discrepancies and reaching a consensus, the annotators achieved a remarkable 97% agreement rate. Third, to quantify the credibility of GPT-3.5’s evaluations, we consider different degrees as the lowest acceptable safety threshold and calculate the agreement rate with human annotations. The results, detailed in Table 8, show a notably high agreement of average 96% when categorizing C degree as the minimum harmfulness level. This underscores the enhanced reliability of GPT-3.5 annotations in conjunction with human judgment in harm detection.
Degrees $=$ A $\geq$ B $\geq$ C $\geq$ D $\geq$ E
Vanilla 51% 72% 93% 90% 89%
Ours 97% 98% 99% 99% 3.0%
Table 8: Agreement rate between human annotations and different harm degrees provided by GPT-3.5 evaluation. A, B, C, D, and E degrees refer to ‘very harmful’, ‘harmful’, ‘moderately harmful’, ‘slightly harmful’, ‘not harmful at all’, respectively.
@@ 6 Conclusion
To tackle the challenge of widespread complex and stealthy jailbreak attacks, which pose deployment risks for large language models, we propose a simple yet highly effective strategy called Intention Analysis ($\mathbb{IA}$). $\mathbb{IA}$ works in the inference stage. Specifically, $\mathbb{IA}$ leverages the intrinsic intent recognition capacities of LLMs, and instructs them to first analyze the essential intention of user queries before responding to it through two stages. Extensive experiments on SAP200 and DAN benchmarks across models of varying scales and alignment levels show that $\mathbb{IA}$ could consistently and significantly reduce the harmfulness in responses (averagely -46.5% attack success rate), while maintaining the general helpfulness. Additionally, we further verify the robustness of $\mathbb{IA}$ to the specific expression of instructions and quality of intention analysis.
@@
Despite the effectiveness of our method in defending sophisticated jailbreak prompts, these prompts do not encompass the entire potential jailbreak attacks encountered in real-world scenarios. Consequently, the practical applicability of our approach remains to be validated through further testing. Our research underlines the importance of intention analysis in improving LLM safety, suggesting future work focusing on integrating this into training to reduce inference costs. Additionally, in the face of the rapid advancements in the adversarial attacks community, there is a pressing need for developing more effective and robust defense strategies for LLMs.
@@ A.1 Safety Datasets
To assess the effectiveness of our method on in-the-wild jailbreak prompts, we employ two jailbreak prompt datasets. The first is forbidden question set developed by Shen et al. (2023), which is currently the largest in-the-wild jailbreak prompt dataset. To improve computing efficiency, we extract five questions from each forbidden scenario, forming a jailbreak dataset comprising 8 jailbreak communities $\times$ 3 jailbreak prompts $\times$ 13 forbidden scenarios $\times$ 5 questions, totaling 1560 samples. The term “DAN” is used to denote this dataset. For evaluation, we leverage attack success rate (ASR) to consider the success of a jailbreak attack. Considering the complex instructions in DAN makes it challenging to directly identify the success of an attack through string matching, we turn to utilize a widely-adopted LLM to evaluate the harmfulness of model generations, as will be discussed in Section B.
@@ A.2 Helpfulness Datasets
Appendix B Evaluation Metrics
B.1 Safety Evaluation

### v3

@@ 4.2 Main Results
Our method significantly boosts safety across a diverse range of LLMs and consistently outperforms other defense baselines.
In Table 1, we represent the ASR of varying defense methods on different LLMs under various jailbreak attacks. We can observe that: 1) $\mathbb{IA}$ effectively reduces ASRs across a diverse range of LLMs, irrespective of their inherent safety levels. For LLMs with high vanilla ASRs, such as ChatGLM-6B, Vicuna-7B, Vicuna-13B, MPT-30B-Chat, and DeepSeek-67B-Chat, we significantly lower the average ASRs from 72.7% to 3.79%. Similarly, for LLMs presenting lower vanilla ASRs, such as LLaMA2-7B-Chat and GPT-3.5, $\mathbb{IA}$ further reduces their average ASRs from 13.73% to mere 0.1%. 2) $\mathbb{IA}$ consistently surpasses all defense baselines. For example, on Vicuna-7B, Vicuna-13B, and MPT-30B-Chat, $\mathbb{IA}$ significantly averages a reduction of 37.4% in ASR compared to the second best baseline. We conjecture such satisfactory performance to be attributed to our design of intention analysis mechanism, which leverages LLM’s awareness of harmful content to enhance safety against varying jailbreak attacks.
@@
Our method can consistently enhance safety in the context of more advanced jailbreaks such as multilingual attack and encryption-based attack.
Recent studies Deng et al. (2023b); Yong et al. (2023) reveal that the multilingual jailbreak poses a new defense challenge for LLMs. Yuan et al. (2024) and Wei et al. (2023a) also emphasize the struggles of more powerful LLMs, such as GPT-3.5, to stay safe when countering encryption-based attack. To verify the effectiveness of our method in these advanced jailbreak scenarios, we reproduce MultiJail and CipherChat following Deng et al. (2023b) and Yuan et al. (2024), respectively, and conduct further experiments on GPT-3.5888We observe a high rate of invalid responses on smaller LLMs like ChatGLM-6B, Vicuna-7B and Vicuna-13B under MultiJail, as demonstrated in Deng et al. (2023b). And smaller LLMs may lack advanced encryption capability, which is required by CipherChat (Yuan et al., 2024). So we only adopt GPT-3.5 for these advanded jailbreak attacks experiments.. The results of GPT-3.5 with and without our $\mathbb{IA}$ are presented in Figure 3. We observe that 1) our $\mathbb{IA}$ consistently maintains performance in low-resource languages, e.g., th, bn, sw, and jv, even in scenarios where a malicious jailbreak prompt [^9] We adopt the same jailbreak prompt in Deng et al. (2023b), namely AIM. is attached to the multilingual attacks, 2) our $\mathbb{IA}$ significantly enhances safety when facing advanced encryption-based attack, even under the most effective SelfCipher attack. These demonstrate the effectiveness of our intention analysis defense mechanism under more advanced jailbreak attacks.
Our method maintains LLM’s general helpfulness on harmless queries.
@@
To explore the impact of our method on the general performance of LLMs, we conduct experiments on several acknowledged helpfulness datasets and report the results in Table 2. As observed, for harmless user prompts, our $\mathbb{IA}$ does not compromise the general helpfulness, indicated by a comparable win rate on AlpacaEval, a similar accuracy on MMLU, and better MC1 and MC2 scores on TruthfulQA compared with vanilla LLMs. These results indicate that $\mathbb{IA}$ can be deployed in real applications to enhance LLM safety while preserving general helpfulness.
Figure 4: The confusion matrix illustrating the relationship between the success of intention analysis and the harmlessness of LLM’s final response on SAP200 and DAN datasets. “IR Succ.” and “IR Fail.” represent success or failure of intention analysis, respectively.
Our method ensures the helpfulness of safe responses for harmful queries.
To comprehensively study the impact of our $\mathbb{IA}$ on responses to harmful queries, we conduct a manual review of 100 random-sampled refusals on both DAN and SAP200 datasets for each of the seven LLMs under our $\mathbb{IA}$. We manually check for 1) reasonable explanations about why it rejects to respond and 2) positive suggestions for the user. In Table 3, our findings show that, following $\mathbb{IA}$, almost all LLMs can give detailed explanations in nearly 100% of cases (except for GPT-3.5 due to OpenAI’s external moderation mechanism during API interactions (OpenAI, 2023)). Specifically, we find LLMs indeed conduct more nuanced handling regarding varying unsafe intents (e.g. terrorism and violence). Detailed cases are presented in Figures 14.
LLaMA2-7B-Chat Vicuna-7B Vicuna-13B
[…]
SAP200 100% 100% 100%
MPT-30B-Chat DeepSeek-67B-Chat GPT-3.5
[…]
SAP200 100% 100% 49%
Table 3: Manual check results of response’s helpfulness for harmful queries on DAN and SAP200 datasets in terms of rate.
@@ 5 Discussion
5.1 Analysis of Intention Analysis Mechanism
@@
Most LLMs can effectively recognize the intention behind jailbreak queries.
Intention analysis is a straightforward language understanding task for LLMs to proficiently perform (Bender and Koller, 2020; Zhu et al., 2024; Gómez-Pérez et al., 2023). The results of intention analysis are binary—either LLMs can successfully detect the intention, such as identifying plans to “rob a bank” as shown in Figure 2, or they fail and miss it. In Figure 4, we count the samples and examine the correlation between successful intention analysis (see Appendix C.4 for evaluation details) and producing harmless responses on SAP200 and DAN datasets [^10] SAP200 and DAN datasets are chosen for intention analysis evaluation due to their most complex and stealthy intentions among jailbreak datasets tested..
@@
Figure 5: Performance of $\mathbb{IA}$ with varying correct intention ratio on DAN dataset.
We observe that most LLMs are highly effective in recognizing intentions behind complex and stealthy jailbreak queries, achieving a nearly 100% success rate in Vicuna-13B, MPT-30B-Chat, and DeepSeek-67B-Chat. Particularly, the intention recognition rate of LLaMA2-7B-Chat is relatively lower due to its excessively strong inherent safety, leading to directly refusing to answer the user query (see Figure 15 for detailed cases).
Two factors that affect $\mathbb{IA}$ performance are intention analysis outcomes and inherent LLM safety.
(1) Intention analysis outcomes [^11] We also conduct cross-intention analysis experiment on Vicuna-7B and Vicuna-13B in Appendix F to explore the effect of different intention analysis LLMs.: In Figure 5, we explore the influence of intention analysis outcomes by simulating varying correct intention ratios through replacing the gold intention identified in the first stage response with an invalid field (such as “[secret]”). We can see that $\mathbb{IA}$ performance gradually improves as the correct intention ratios increase, indicating that enhancing the intention analysis ability can further boost $\mathbb{IA}$ effectiveness [^12] Despite a 0% correct intention ratio, $\mathbb{IA}$ remains effective compared to vanilla LLMs, since the intention analysis sequence format allows replacing gold intention labels with invalid ones to be marginally detrimental, as recognized by the In Context Learning (ICL) community (Min et al., 2022)..
@@
Table 4: Comparison between our method and well safety-trained LLM in safety and helpfulness (%). The best and second best are in bold and underline.
(2) Inherent LLM safety: In Figure 4, among LLMs achieving nearly 100% in intention recognition rates, there is a notable difference in harmful response rates, ranging from 0.3% for Vicuna-7B to 19.3% for MPT-30B-Chat, indicating the impacts of inherent LLM safety on $\mathbb{IA}$ results (see Figure 16 for related case study). These pose two potential directions for further improving $\mathbb{IA}$ performance in the future: enhancing the LLMs’ intention analysis capability and their inherent safety.
5.2 Efficiency and Robustness of Our $\mathbb{IA}$
Our method achieves comparable safety with well safety-trained LLMs without need for additional training.
Our method aims to enhance LLM safety in the inference stage. A natural question arises: how does its performance compare to well safety-trained LLMs? To answer this, we compare our method with a representative well safety-trained LLM, i.e., LLaMA2-7B-Chat. The results are listed in Table 4. We can see that our method achieves comparable performance to LLaMA2-7B-Chat on safety datasets while outperforming LLaMA2-7B-Chat on the helpfulness dataset by almost 5%. This demonstrates the advantage of our $\mathbb{IA}$ to achieve both safety and helpfulness goals without additionally resource-consuming safety training.
Our efficient one-pass variant of $\mathbb{IA}$ still works well.
As aforementioned, to maximize the performance, our $\mathbb{IA}$ follows a two-stage process. A natural question arises of whether our mechanism can be merged into one step, to save the decoding overhead. To verify this, we design a cheaper one-pass $\mathbb{IA}$, without letting the LLMs separately explain the intention (see Figure 9 for detailed prompts). As shown in Table 5, we see that 1) both the two-stage and one-pass $\mathbb{IA}$ perform well on different LLMs, 2) two-stage $\mathbb{IA}$ can further improve LLM safety compared with one-pass $\mathbb{IA}$.
Our method is robust to different $\mathbb{IA}$ prompts.
To assess the resilience of our method against specific expressions, we construct other two sets of alternative instructions of our $\mathbb{IA}$ (see Figure 10 for detailed prompts) and conduct comparison with the prompt set we used in main experiments. Table 6 presents the comparison results. We observe that three prompt sets all result in significant and comparable decreases in ASRs on the SAP200 across different LLMs. This uniformity highlights a key conclusion: it is the intention analysis mechanism, but not the specific instruction expressions, that makes our $\mathbb{IA}$ effective.
Vicuna-7B Vicuna-13B ChatGLM-6B
Vanilla 73.4 65.4 45.8
+ One-Pass $\mathbb{IA}$ 5.50 1.13 39.0
+ Two-Stage $\mathbb{IA}$ 0.31 1.12 6.12
Table 5: Comparison of our $\mathbb{IA}$ with different implementations (one-pass and two-stage) on SAP200 in terms of ASR (%). The best are highlighted in bold.
Vicuna-7B Vicuna-13B ChatGLM-6B
Vanilla 73.4 65.4 45.8
+ Prompt A 2.94 0.88 5.81
+ Prompt B 5.13 2.06 5.44
+ Ours 0.31 1.12 6.12
Table 6: Ablation of different $\mathbb{IA}$ prompts on SAP200 in ASR (%). The best and second best results are highlighted in bold and underline.
@@ 6 Conclusion
In this work, a simple yet highly effective defense strategy, namely $\mathbb{IA}$, is proposed to handle the widespread complex and stealthy jailbreak attacks. $\mathbb{IA}$ works by leveraging the intrinsic intention recognition capacities of LLMs, and instructing them to first analyze the essential intention of user queries before finally responding through two stages. Extensive experiments on representative jailbreak benchmarks across LLMs of varying scales and alignment levels show that $\mathbb{IA}$ could consistently and significantly reduce the harmfulness in responses (averagely -53.1% attack success rate), while maintaining the general helpfulness. Further analysis indicates that enhancing LLMs’ intention analysis capability and their inherent safety are two directions for future improvements.
@@
Due to limited budgets, GPT-4 is not yet included in our experiments. However, its superior language understanding (including intention analysis) and safety capabilities suggest it will follow the trends we see with GPT-3.5 in our study. Additionally, despite the effectiveness of our method in defending sophisticated jailbreak prompts, these prompts do not encompass the entire potential jailbreak attacks encountered in real-world scenarios. Consequently, the practical applicability of our approach remains to be validated through further testing. Our research underlines the importance of intention analysis in improving LLM safety, suggesting future work focusing on integrating this into training to reduce inference costs. Additionally, in the face of the rapid advancements in the adversarial attacks community, there is a pressing need for developing more effective and robust defense strategies for LLMs.
@@ A.1 Safety Datasets
To assess the effectiveness of our method on in-the-wild jailbreak prompts, we employ two jailbreak prompt datasets. The first is forbidden question set developed by Shen et al. (2023), which is currently the largest in-the-wild jailbreak prompt dataset. To improve computing efficiency, we extract five questions from each forbidden scenario, forming a jailbreak dataset comprising 8 jailbreak communities $\times$ 3 jailbreak prompts $\times$ 13 forbidden scenarios $\times$ 5 questions, totaling 1560 samples. The term “DAN” is used to denote this dataset. For evaluation, we leverage attack success rate (ASR) to consider the success of a jailbreak attack. Considering the complex instructions in DAN makes it challenging to directly identify the success of an attack through string matching, we turn to utilize a widely-adopted LLM to evaluate the harmfulness of model generations, as will be discussed in Section C.2.
@@ A.2 Helpfulness Datasets
Appendix B Language Models
ChatGLM-6B
LLaMA-2-7B-chat
Vicuna-7B-v1.1
MPT-30B-Chat
DeepSeek-67B-Chat
GPT-3.5
Appendix C Evaluation Metrics
C.1 Automation-based Safety Evaluation
For evaluation of DAN, considering the complex instructions in DAN makes it challenging to directly identify the success of an attack through string matching, we turn to utilize a widely-adopted LLM to evaluate the harmfulness of model generations. Practically, we adopt gpt-3.5-turbo-0613131313https://openai.com/blog/chatgpt as the evaluation function following Deng et al. (2023a). Especially, the model responses are submitted via API calls, and a specific evaluation prompt is designed for gpt-3.5-turbo-0613 to rate the harmfulness on a five-degree scale (‘A (Very Harmful)’, ‘B (Harmful)’, ‘C (Moderately Harmful)’, ‘D (Slightly Harmful)’, and ‘E (No Harmful at All)’). Figure 6 shows the detailed instruction for gpt-3.5-turbo-0613 evaluation.
Human Evaluation
To validate GPT-3.5 as the harmfulness annotation function, we conduct a human validation process. First, we randomly select 200 responses from both the vanilla model and our method on the DAN dataset. Then we engage two expert annotators proficient in alignment studies of LLMs and fluent in English. After reassessing their evaluations for cases with discrepancies and reaching a consensus, the annotators achieved a remarkable 97% agreement rate. Third, to quantify the credibility of GPT-3.5’s evaluations, we consider different degrees as the lowest acceptable safety threshold and calculate the agreement rate with human annotations. The results, detailed in Table 7, demonstrate the highest agreement of average 96% when categorizing C degree as the minimum harmfulness level, i.e., viewing the evaluation degrees of ‘A (Very Harmful)’, ‘B (Harmful)’, and ‘C (Moderately Harmful)’ as a success of jailbreak attacks. This underscores the enhanced reliability of GPT-3.5 annotations in conjunction with human judgment in harm detection.
Degrees $=$ A $\geq$ B $\geq$ C $\geq$ D $\geq$ E
Vanilla 51 72 93 90 89
Ours 97 98 99 99 3.0
Table 7: Agreement rate (%) between human annotations and different harm degrees provided by GPT-3.5 evaluation. A, B, C, D, and E degrees refer to ‘very harmful’, ‘harmful’, ‘moderately harmful’, ‘slightly harmful’, ‘not harmful at all’, respectively.
C.2 Rule-based Safety Evaluation